### v1

# SAM2-ELNet: Label Enhancement and Automatic Annotation for Remote Sensing Segmentation

Abstract
Remote sensing image segmentation is crucial for environmental monitoring, disaster assessment, and resource management, directly affecting the accuracy and efficiency of surface information extraction. The performance of existing supervised models in remote sensing image segmentation tasks highly depends on the quality of label data. However, current label data mainly relies on manual annotation, which comes with high time costs and is subject to subjective interference, resulting in distortion of label boundaries and often a loss of detail. To solve the above problems, our work proposes an Edge-enhanced Labeling Network, called SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism. This model effectively addresses issues such as label detail loss, fragmentation, and inaccurate boundaries. Due to the scarcity of manually annotated remote sensing data, the feature extraction capabilities of traditional neural networks are limited. Our method uses the Hiera backbone of the pre-trained self-supervised large model segment anything model 2 (SAM2) as the encoder, achieves high-quality and efficient feature extraction even with small samples by fine-tuning on downstream tasks. This study compared the training effects of original and enhanced labels on the manually annotated Deep-SAR Oil Spill (SOS) dataset. Results showed that the model trained with enhanced labels performed better and had a lower final loss, indicating closer alignment with the real data distribution. Our work also explores the potential of extending the model into an efficient automatic annotation framework through generalization experiments, facilitating large-scale remote sensing image interpretation and intelligent recognition.
Index Terms:
Remote Sensing Image Segmentation, Automatic Annotation, Edge Attention Mechanism, Binarization Processing, Label Enhancement.
I Introduction
Remote sensing typically refers to the technology of imaging the earth system in a specific electromagnetic spectrum band on an aerospace or aviation platform, so as to obtain multi-faceted characteristic information of the observed object. It’s imaging methods [1] include optical imaging, thermal infrared [2], hyperspectral [3], [4], and synthetic aperture radar (SAR) [5]. Currently, there are already various ways to extract valuable information from the massive amount of data. Remote sensing image segmentation has become a focal point of contemporary research due to its capability to delineate target regions and reveal terrestrial spatial structures.
Remote sensing image segmentation methods are mainly divided into traditional methods and deep learning-based methods. Traditional segmentation methods typically employ techniques such as thresholding, random forest [37], support vector machine [38], [16], and conditional random field [17] to delineate different surface structure regions. When applied to simple scenes, these methods often yield satisfactory results. But their sensitivity to radiometric noise, occlusive shadows, and topological ambiguities leads to issues, especially when dealing with high-dimensional features and complex spatial relationships.
Driven by advancements in the field of computer vision (CV), deep learning techniques have become the leading approach for remote sensing image segmentation [49], [50], [51], [52], [53]. Supervised learning methods in this domain rely on large-scale, high-quality annotated datasets to effectively train robust models. However, most existing datasets are manually annotated. The inherent heterogeneity of targets in remote sensing images, coupled with noise interference, affects boundary clarity, and increases the difficulty of precise recognition. Due to the subjectivity of annotators and the inherent complexity of the task, annotators often generate annotations of suboptimal quality during the labeling process, which tend to omit details or over-simplify geometric shapes[6]. As dataset scales expand exponentially, the corresponding annotation costs grow prohibitively. Therefore, reducing reliance on high-quality manual annotations, improving annotation efficiency and accuracy, and developing effective automated annotation frameworks are crucial for advancing intelligent remote sensing recognition and interpretation.
Interestingly, the computer vision community has encountered challenges similar to those in remote sensing. Research efforts are increasingly converging on strategies to alleviate the dependency on manually annotated training samples. self-supervised learning (SSL) extracts meaningful patterns from large-scale unlabeled data, reducing annotation costs. Dosovitskiy et al. [7] introduced the vision transformer (ViT), which brings the transformer architecture designed for natural language processing into visual tasks, leading to more efficient feature encoding. Meanwhile, contrastive learning [9], exemplified by methods like simple framework for contrastive learning of visual representations (SimCLR) [11] and momentum contrast for unsupervised visual representation learning (MoCo) [10], refines feature representation by distinguishing between positive and negative sample pairs, leading to more stable and generalizable embeddings. Unlike contrastive learning methods, which rely on explicit positive and negative sample pairs, Grill et al. [54] proposed bootstrap your own latent (BYOL), which adopts a momentum-based update strategy for the Target Network without requiring negative samples. This enables contrast-free self-supervised learning and leads to more robust feature representations. In recent years, the emergence of large foundation models in computer vision has provided stronger support for downstream tasks. Caron et al. [8] introduced self-distillation with no labels (DINO), which leverages self-distillation with a student-teacher network to learn powerful visual representations without requiring labeled data. He et al. [12] proposed masked autoencoders (MAE), which enhance visual representations by randomly masking image regions and training a decoder for reconstruction. Building on MAE’s success, the segment anything model (SAM) [13] and its enhanced version, SAM2 [14], are pretrained on larger-scale datasets, endowing SAM2 with outstanding feature extraction capabilities.
With the growing adoption of these methods, researchers are seeking better ways to optimize manually annotated data while improving labeling efficiency in remote sensing applications. Compared with CV tasks, remote sensing images have more severe noise and scale changes, which makes end-to-end self-supervised methods face greater challenges. Wang et al. [60] proposed an automatic annotation tool that integrates multiple deep learning models (including SAM), called X-Anylabeling. Its application to remote sensing image annotation still requires manual input, and its effectiveness remains limited in scenarios involving small object detection, high-noise images, or abnormal data distributions. To enhance the application of SAM in remote sensing image segmentation, Zhang et al. [36] introduced the adapter-scale and adapter-feature modules to improve SAM’s encoding capability, making it more suitable for semantic segmentation tasks in remote sensing. Ma et al. [35] proposed the SGO-SGB framework by incorporating object consistency loss and boundary preservation loss, addressing SAM’s fragmentation and boundary inaccuracy issues in semantic segmentation of remote sensing imagery. Chen et al. [61] developed the RSPrompter method, which learns to generate category-specific prompts to enhance SAM’s applicability in remote sensing instance segmentation. Despite SAM2 performs well in many tasks, it may still generate segmentation results unrelated to specific categories when manual prompts are absent, which limits its usability in the specific context of remote sensing segmentation.
To address this limitation, Xiong et al. [15] integrated SAM2’s hierarchical feature extraction with U-Net’s spatial attention mechanism and proposed a novel network, SAM2-UNet. The architecture combines U-Net’s symmetrical decoder with a lightweight parameter adapter [62], [63], enabling efficient fine-tuning for downstream applications. Extending the SAM2-UNet framework, we propose a new model for label refinement and automatic annotation, called edge-enhanced labeling network (SAM2-ELNet). This model leverages large-scale self-supervised pretrained models for feature extraction and integrates few-shot learning for fine-tuning on downstream tasks. Our work introduce an edge attention mechanism and a labeling module to address issues such as label detail loss, fragmentation, and inaccurate boundaries in manually annotated remote sensing segmentation data, with the potential to be extended into an efficient automatic labeling framework.The main contributions of this study are as follows.
  1. 1.
    To address the issue of suboptimal annotations caused by heterogeneity and noise interference in remote sensing segmentation tasks, our work propose SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism. This model effectively optimizes issues such as label detail loss, fragmentation, and inaccurate boundaries.
  2. 2.
    To enhance the applicability and training efficiency of SAM for specific remote sensing image segmentation tasks, we leverage a pre-trained large model and adopt a fine-tuning strategy with a frozen encoder. This approach retains the robust feature extraction capability of SAM2 while significantly improving training efficiency.
  3. 3.
    In exploring large-scale automatic annotation, we discuss the model’s generalization ability and investigate its potential to be extended into an efficient automatic labeling framework, facilitating the construction of scalable, high-quality datasets.
The remainder of this paper is organized as follows. Section II reviews related works on remote sensing image segmentation, automatic annotation methods, and the application of pretrained models in remote sensing. Section III presents the proposed SAM2-ELNet architecture, providing a detailed description of its components. Section IV offers a comprehensive experimental analysis, including model performance evaluation, ablation studies, generalization capability assessment, and efficiency comparisons. Both qualitative and quantitative analyses are conducted to validate the effectiveness of the proposed model. Finally, Section V concludes the study.
II Related Works
II-A Advances in Annotation Strategies for Remote Sensing Image Segmentation
In recent years, researchers have actively explored efficient annotation and training strategies to reduce dependence on high-quality manual annotations, lower costs, and enhance the accuracy of remote sensing image segmentation.
Yang et al.[18] introduced the EasySeg framework, which addresses domain adaptation challenges in remote sensing semantic segmentation. Their method specifically targets domain discrepancies and misperception issues in cross-domain applications. It incorporates a ’See-First-Ask-Later’ point-level annotation strategy and an Interactive Semantic Segmentation Network (ISS-Net), substantially reducing annotation costs while maintaining segmentation accuracy under weak supervision. Liu et al.[19] examined the role of noisy labels in supervised pretraining, analyzing their impact on encoder performance in remote sensing segmentation tasks. By comparing noisy label pretraining with self-supervised learning methods such as DINO and MoCo, they further demonstrated the effectiveness of noisy labels in pretraining and examined how class inconsistency influences model transferability. Li et al.[20] developed the SAM-OIL framework, which integrates YOLOv8-based object detection, an adapted SAM segmentation module, and an ordered mask fusion (OMF) mechanism for oil spill detection in SAR imagery. This method significantly improves detection accuracy while reducing dependence on fine-grained annotations.
II-B Enhancing Annotation Quality and Efficiency with Advanced Learning Models
To further enhance annotation quality and efficiency, Wu et al.[21] pioneered the use of vision-language models (VLM) in farmland segmentation. They constructed a farmland image-text paired dataset and integrated semantic segmentation models with large language models (LLMs), developing the FSVLM model, which improved generalization and segmentation accuracy of remote sensing segmentation . Li et al.[22] introduced a semi-automatic annotation method that combines anchor real-time lines and ROI real-time line algorithms to address label incompleteness and deficiencies in traditional remote sensing indices and interactive segmentation. Han et al.[23] designed the DM-ProST framework, which generates high-quality pseudo-labels and expands training datasets through mutual correction between two deep learning models. This approach selects reliable unlabeled samples, improving segmentation accuracy. Zhu et al.[24] proposed AIO2 (adaptively triggered online object label correction), which integrates adaptive correction triggering (ACT) and online object label correction (O2C) to enhance robustness against noisy labels. Shi et al.[25] developed an automated multi-temporal change detection method for remote sensing images, incorporating boundary preservation techniques. By applying adaptive cropping and optimized classifier training, their method generates high-quality annotations and updates existing polygon-based labels. However, it does not resolve label omissions and struggles with non-polygonal targets such as oil spills and lakes.
II-C Optimization and Application of SAM in Remote Sensing and Geoscience
The segment anything model has gained widespread attention for its role in self-supervised foundation models across remote sensing, medical imaging, and geophysical exploration[26]. In land use and crop type segmentation, SAM significantly improves performance[27], [28]. Ma et al.[29] explored geophysical foundation models (GeoFM) for geophysical exploration and applied SAM to seismic phase picking. SAM also enhances building segmentation and urban area delineation in remote sensing imagery [30], [31], [32], [33], [34], demonstrating its adaptability in processing complex geographical features.
Figure 1: Architecture of SAM2-ELNet. The proposed SAM2-ELNet comprises an encoder, decoder, receptive field blocks (RFB), and a labeling module. It integrates an edge attention mechanism and parameter-efficient fine-tuning (PEFT) to enhance segmentation accuracy and improve label quality for remote sensing imagery.
III Proposed Method
Building upon the research advancements of SAM2-Unet [15], this paper proposes a model SAM2-ELNet designed for enhanced labeling and automatic annotation. This model is specifically tailored for annotation enhancement and automatic labeling, integrating adapters to achieve parameter-efficient fine-tuning (PEFT) while preserving the feature extraction capability of the pre-trained encoder.
As shown in Fig.1, Since remote sensing datasets are limited, retraining Hiera from scratch could weaken its feature extraction ability. To preserve its effectiveness, we adopt a frozen backbone strategy during training. Remote sensing images exhibit higher levels of noise and heterogeneity. To enable efficient fine-tuning while maintaining adaptability to remote sensing tasks, SAM2-ELNet inherits the adapter design from SAM2-Unet. We integrate adapters before the multi-scale modules of the Hiera encoder, allowing the model to dynamically recalibrate feature distributions and reducing the number of tunable parameters. SAM2-ELNet employs a frozen encoder strategy, its PEFT approach is similar to low-rank adaptation [39] (LoRA). LoRA mainly targets the attention mechanism of Transformer structures, but SAM2-ELNet embeds its adapter module before the multi-scale modules of the Hiera encoder to adjust feature mapping and enhance model adaptability, making it more suitable for remote sensing label optimization tasks.
In the decoder stage, SAM2-ELNet incorporates an edge attention mechanism to sharpen feature responses along object boundaries, improving segmentation accuracy. We also introduce a labeling module that transforms continuous grayscale maps into high-fidelity binary labels, preserving critical details. The subsequent sections will delve deeper into each module.
III-A Encoder, Adapter, and RFB Modules
Hiera (hierarchical transformer) [40] is a Transformer backbone network designed for efficient visual feature extraction. Unlike traditional convolutional neural networks (CNNs), Hiera captures both local details and global semantics through hierarchical feature modeling. Its layered computation strategy reduces computational cost while maintaining strong representational power. Our model uses the SAM2-pretrained Hiera as an encoder, utilizing its multi-scale feature learning to generate high-quality segmentation features. Four receptive field blocks (RFBs) [55], [56] further process the extracted features, reducing the channel count to 64 to enhance lightweight feature representation.The structure of the Hiera is illustrated in Fig. 2.
Hiera’s large parameter count makes full fine-tuning computationally expensive. To enhance parameter efficiency, we adopt the SAM2-Unet design, freezing Hiera’s parameters and inserting lightweight adapters before each multi-scale block. Each adapter includes a downsampling linear layer followed by a GeLU [57] activation, then an upsampling linear layer with another GeLU activation. This structure allows the model to efficiently recalibrate feature distributions while reducing the number of trainable parameters.
III-B Decoder
The original decoder follows the classical U-Net design, consisting of three decoder blocks, each comprising two Conv-BN-ReLU layers. When processing remote sensing images with blurred boundaries or low contrast, this structure tends to produce segmentation outputs with unclear contours, fragmented edges, or excessive smoothing. To address these issues, we introduce the edge attention mechanism (EAM) to enhance segmentation precision, particularly in boundary regions.
Figure 2: Architecture of the Hiera. Proposed in SAM2-Unet, the encoder leverages the Hiera backbone for multi-scale feature modeling, with adapters enhancing feature adaptation.
EAM enhances the feature map’s focus on boundary regions during the upsampling stage. After each upsampling step, it ensures that multi-scale feature aggregation preserves critical boundary information. The EAM in this study employs convolution operations to extract object boundary features and applies element-wise weighting to refine the input features, guiding the model to emphasize boundary details. The edge attention weight is computed as follows:
$A_{E}=\sigma(\text{BN}(W_{\text{conv}}*X))$ (1)
where $X$ is the input feature map, $W_{\text{conv}}$ is a 3×3 convolution kernel. * denotes the convolution operation, BN refers to batch normalization, $\sigma$ denotes the Sigmoid function.
The generated edge attention weight $A_{E}$ modulates the original feature $X$ to achieve edge enhancement:
$X^{\prime}=X\odot A_{E}$ (2)
where $\odot$ represents element-wise multiplication. $X^{\prime}$ is an enhanced feature map emphasizing edge-related information. This module adopts a lightweight design, using only a independent 3×3 convolutional layer, which significantly reduces computational cost compared to self-attention mechanisms.
III-C Loss Function
We formulate the loss function as a combination of the weighted Intersection over Union (IoU) loss $L_{\text{wIoU}}$ and the weighted binary cross-entropy (BCE) loss $L_{\text{wBCE}}$, defined as:
$L=L_{\text{wIoU}}+L_{\text{wBCE}}$ (3)
To enhance training effectiveness, we apply deep supervision to all segmentation outputs $S_{i}$, leading to the final total loss function:
$L_{\text{total}}=\sum_{i=1}^{3}L(G,S_{i})$ (4)
where $G$ denotes the ground truth labels, and $S_{i}$ represents the multi-level segmentation outputs generated by the model.
Figure 3: Labeling Module Architecture. Converts grayscale predictions into binarized labels using edge detection, morphological processing, and the watershed algorithm for precise region extraction.
III-D Labeling Module
The model’s initial predictions are grayscale images with intensity values ranging from 0 to 255. To transform these predictions into binarized labels, our work design a labeling module based on edge detection and the watershed transform. This approach integrates Canny edge detection [41], morphological processing [42], [43], distance transform, and the watershed algorithm [44], [45] to accurately extract foreground, background, and boundary regions. The structure of the labeling process is illustrated in Fig. 3.
To detail the labeling procedure, we apply the Canny operator to an input image $I$ to extract object contours, emphasizing prominent edges. This process produces the edge map $E$, defined as:
$E=\text{Canny}(I,T_{\min},T_{\max})$ (5)
where $E$ represents the binary edge map, while $T_{\min}$ and $T_{\max}$ denote the lower and upper threshold values of the Canny operator. This step enhances object boundaries by highlighting regions with sharp intensity transitions.
To fill gaps within foreground objects, we apply a combination of binarization and morphological dilation. After thresholding the input image, we perform dilation:
$T(x,y)=\begin{cases}255,&I(x,y)<T_{\text{thresh}}\\ 0,&\text{otherwise}\end{cases}$ (6)
$D=\text{dilate}(T,K,n)$ (7)
where $T_{\text{thresh}}$ is the binarization threshold, $K$ is the morphological kernel, and $n$ represents the number of dilation iterations. This operation enhances the foreground region, refines boundary continuity, and reduces discontinuities within the object.
To generate the segmentation labels, we construct the foreground mask $F_{\text{fg}}$ by merging the dilation result $D$ with the Canny edge map $E$, Similarly, we obtain the background mask $F_{\text{bg}}$ by inverting the binarization result:
$F_{\text{fg}}=D\cup E,\quad F_{\text{bg}}=\neg T$ (8)
Next, we optimize the foreground segmentation by constraining it with the Euclidean distance transform:
$D_{T}=\text{distanceTransform}(F_{\text{fg}},L_{2},5)$ (9)
$F_{\text{fg}}^{\prime}=\begin{cases}255,&D_{T}(x,y)>\alpha\cdot\max(D_{T})\\ 0,&\text{otherwise}\end{cases}$ (10)
where $\alpha$ is the threshold scaling factor, which controls the extent of the foreground region to ensure boundary precision without excessive expansion.
Using the extracted foreground and background regions, the labeling matrix $M(x,y)$ is constructed as:
$M(x,y)=\begin{cases}2,&(x,y)\in F_{\text{fg}}^{\prime}\\ 1,&(x,y)\in F_{\text{bg}}\end{cases}$ (11)
This matrix serves as the marker input for the watershed transform, which segments the image into foreground, background, and boundary regions based on gradient-based region growing:
$M^{\prime}=\text{Watershed}(I,M)$ (12)
The final labeled output $M^{\prime}$ assigns 0 for boundary regions and background, sets 255 for foreground objects, thereby generating high-quality segmentation labels.
III-E Overview of the Label Enhancement Process
The overall workflow of the proposed label enhancement method consists of the following stages:
  1. 1.
    Initial Manual Annotation. A small subset of manually annotated data serves as the initial training set. Although these annotations may exhibit suboptimal quality, they still provide essential supervisory signals. This dataset supports the preliminary training of the model, enabling it to learn fundamental segmentation patterns and adapt to the specific characteristics of the target task.
  2. 2.
    Model Fine-Tuning. The model initializes with pretrained weights and fine-tunes on the manually annotated dataset. With the encoder weights remaining frozen, the fine-tuning process primarily optimizes the adapters and decoder. This strategy mitigates the adverse impact of suboptimal quality annotations on model learning, facilitating better adaptation to remote sensing segmentation tasks.
  3. 3.
    Label Generation. Upon completion of training, the fine-tuned model infers segmentation results, automatically generating high-quality segmentation labels.
  4. 4.
    Iterative Optimization (Optional). The model can further fine-tune using the enhanced labels, iteratively refining segmentation quality through adaptive learning. This iterative strategy progressively improves annotation accuracy and stability.
IV Experiments and Discussion
IV-A Datasets
The Deep-SAR Oil Spill (SOS) dataset consists of SAR images from two sources [46]: PALSAR data from the Mexican oil spill region and Sentinel data from the Persian Gulf. PALSAR, an L-band SAR sensor on the ALOS satellite, and Sentinel, a C-band SAR sensor, both provide all-weather, cloud-free imaging. In this dataset, dark regions in SAR images typically indicate oil spill contamination. Zhu et al.[46] constructed the dataset from 21 SAR images, expanding it through cropping, rotation, and noise addition to produce 6,456 oil spill images $416\times 416$. The Mexican oil spill subset includes 3,101 training and 776 test images, while the Persian Gulf subset contains 3,354 training and 839 test images. Manual interpretation and GIS expert sampling generated the annotations, with optimizations addressing class imbalance.
Ideally, oil spill regions should have smooth and continuous boundaries. Due to the inherent subjectivity of manual annotation, as well as the heterogeneity of remote sensing images and noise interference, annotation results often exhibit issues such as missing label details, fragmentation, and polygonal boundaries. Illustrative examples are shown in Fig. 4.
(a) Test sample 10011
(b) Test sample 10074
(c) Train sample 10779
(d) Train sample 11929
Figure 4: Examples of annotation inconsistencies in the SOS dataset. (a) The annotation erroneously classifies a sea region as an oil spill. (b) The annotated boundary exhibits an irregular polygonal shape, deviating from the original SAR image. (c) The label incorrectly expands a small oil spill into a large irregular polygon, while a portion of the oil spill on the left remains unlabeled. (d) The annotation misclassifies a sea region as an oil spill and exhibits irregular polygonal boundaries.
IV-B Hyperparameter Settings
The experiment runs on an NVIDIA GeForce RTX 4090 GPU with 24GB of memory and utilizes CUDA 12.4 for accelerated computation. The model fine-tunes using the pretrained weights from sam2_hiera_large and optimizes with the Adam optimizer. The hyperparameters are set as follows: 200 training epochs, a batch size of 12, a learning rate of 0.001 and a weight decay of $5\times 10^{-4}$.
IV-C Evaluation Metrics
To assess the quality of segmentation results, we use Accuracy and Mean Intersection over Union (MIoU) as evaluation metrics.
IV-C1 Overall Accuracy (ACC)
Overall accuracy measures the model’s ability to correctly classify all categories. It is defined as:
$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}$ (13)
where TP denotes the number of correctly predicted foreground pixels, TN represents the number of correctly predicted background pixels, FP refers to the number of pixels incorrectly classified as foreground, and FN indicates the number of pixels incorrectly classified as background.
IV-C2 Mean Intersection over Union (mIoU)
Mean Intersection over Union (mIoU) evaluates the overlap between the predicted segmentation and the ground truth. It is computed as:
$\text{mIoU}=\frac{1}{N}\sum_{c=1}^{N}\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}+\text{FN}_{c}}$ (14)
where TPc, FPc, and FNc represent the true positive, false positive, false negative counts for class $c$, and $N$ is the total number of classes.
IV-D Comparative Analysis of Label Enhancement
IV-D1 Qualitative Analysis
The model fine-tunes using only the training subset of the PALSAR dataset for 200 epochs. After training, the model enhances both the training and test set labels of the PALSAR dataset, generating high-quality annotations for subsequent analysis.
From a qualitative perspective, the enhanced labels show substantial visual improvements over the original annotations, as illustrated in Fig. 5.
(a) original image
(b) manual annotation
(c) enhanced label
(d) original image
(e) manual annotation
(f) enhanced label
Figure 5: Results of the label enhancement experiment. Each triplet consists of an original SAR image, a manually annotated label, and a model-enhanced label. The enhanced labels exhibit smoother and more continuous boundaries, improved annotation details, and refined structural consistency.
As depicted in Fig. 5, the enhanced labels exhibit smoother and more continuous boundaries, significantly refining annotation details compared to the original labels. This improvement is not limited to isolated cases but is a consistent trend across the dataset. The enhanced labels demonstrate superior performance in boundary smoothness, object shape precision, and the accurate depiction of complex regions and ambiguous edges, enabling a more precise delineation of target contours. The increased consistency and accuracy of the annotations validate the model’s ability to effectively optimize manually labeled data, offering a more reliable supervisory dataset for remote sensing image segmentation tasks.
IV-D2 Quantitative Analysis
In this experiment, we conduct comparative studies using the open-source implementation of DeepLabv3+ [46], [47] to analyze the impact of annotation quality on model performance. Specifically, we compare the performance of models trained on original labels versus those trained on enhanced labels. All experiments follow identical training conditions, utilizing DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as baseline models. Each model is initialized with the same pretrained weights and trained with a consistent set of hyperparameters for 20 epochs.
TABLE I: Impact of Different Label Qualities on Model Performance (DeepLabv3+ Training Comparison)
Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
MobileNet Original Label 0.9251 0.7436 0.0773
Enhanced Label 0.9523 0.8164 0.0452
ResNet50 Original Label 0.9195 0.7396 0.0815
Enhanced Label 0.9476 0.8092 0.0458
ResNet101 Original Label 0.9210 0.7652 0.0792
Enhanced Label 0.9337 0.8136 0.0453
Experimental results Table I and Fig. 6 indicate that models trained with enhanced labels consistently outperform those trained with original labels in terms of mean Intersection over Union (mIoU) and overall accuracy (ACC). This demonstrates that enhanced labels facilitate more effective learning of representative features. Furthermore, models trained with enhanced labels converge faster and achieve lower final loss values, confirming that enhanced labels more accurately capture the true data distribution. As a result, leveraging enhanced labels significantly enhances overall performance in segmentation tasks.
(a) overall loss curve
(b) detailed loss segment
(c) overall accuracy curve
(d) overall mIoU curve
(e) initial loss curve
Figure 6: Training Convergence and Performance Metrics of DeepLabV3+ ResNet50 on Original and Enhanced Labels. (Loss, accuracy (ACC), and mean Intersection over Union (mIoU) trends during training on original and enhanced labels. Each epoch consists of 190 iterations, resulting in a total of 3,800 iterations over 20 epochs.)
IV-E Generalization Performance Evaluation
To evaluate the model’s generalization ability, we fine-tuned it solely on the PALSAR dataset’s training set, optimizing it over 200 epochs. After training, we used the fine-tuned model to refine both the training and testing labels of the Sentinel dataset, producing higher-quality annotations. This approach helped us assess the model’s adaptability to different data distributions and explore its potential for automatic annotation tasks.
IV-E1 Qualitative Analysis
To visually assess the impact of label enhancement on segmentation performance in the Sentinel dataset, we conducted a comparative analysis through visual inspection. As shown in Fig. 7, models trained with enhanced labels outperform those trained on original labels, demonstrating sharper boundaries, finer segmentation details, and better preservation of target regions. The enhancement process effectively mitigates issues like boundary ambiguity and object omission, resulting in more precise segmentation outcomes.
(a) original image
(b) manual annotation
(c) enhanced label
(d) original image
(e) manual annotation
(f) enhanced label
Figure 7: Generalization Performance of Label Enhancement on the Sentinel Dataset. The enhanced labels improve boundary sharpness and segmentation accuracy, demonstrating the model’s adaptability to different data distributions.
IV-E2 Quantitative Analysis
To further quantify the impact of label enhancement on the model’s generalization performance, we conducted comparative experiments under identical training conditions. Specifically, we used DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as benchmark models. We initialized all models with the same pre-trained weights, applied identical hyperparameter configurations, and trained them for 20 epochs to evaluate their adaptability and generalization across different datasets.
TABLE II: Impact of Different Label Qualities on Model Generalization Performance (Cross-Dataset Transfer Experiment)
Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
MobileNet Original Label 0.8347 0.7024 0.1073
Enhanced Label 0.8941 0.7953 0.0606
ResNet50 Original Label 0.8702 0.7525 0.0976
Enhanced Label 0.9056 0.8166 0.0597
ResNet101 Original Label 0.8491 0.7192 0.0983
Enhanced Label 0.9100 0.8261 0.0554
Experimental results in Table II indicate that although overall performance declined after cross-dataset transfer, the models still exhibited strong automatic annotation capabilities, further confirming the feasibility of the proposed approach for automatic annotation in remote sensing imagery.
IV-F Ablation Study: Effect of the Edge Attention Mechanism
To investigate the impact of the edge attention mechanism on remote sensing image segmentation performance, our work conducted an ablation study. By comparing model configurations with and without the EAM, our work evaluated differences in both qualitative visualization and quantitative metrics to assess its contribution to improving segmentation accuracy.
IV-F1 Qualitative Analysis
Fig.8 presents the segmentation results obtained with (left) and without (right) the Edge Attention Mechanism.
Figure 8: Impact of the Edge Attention Mechanism on Segmentation Results. (b) Without Edge Attention Module, (c) With Edge Attention Mechanism; (e) Without Edge Attention Mechanism, (f) With Edge Attention Mechanism.
Experimental results show that the Edge Attention Mechanism significantly improves segmentation accuracy along object boundaries. In regions with low contrast or blurred edges, this module captures finer details, generating smoother and more continuous segmentation boundaries while reducing jagged or fragmented edges. In contrast, models without the EAM perform worse in these areas, often distorting shapes and creating ambiguous boundaries, which lowers segmentation quality.
However, introducing the EAM may lead to minor information loss in small-scale regions, causing some small oil spill areas to be overlooked. From an annotation perspective, incorporating edge attention is more suitable for generating segmentation labels, especially in scenarios where clear boundary delineation is crucial.
Figure 9: Comparison of Different Labeling Methods. (a) Original SAR Image, (b) Decoder Output, (c) Threshold-Based Binarization, and (d) Proposed Labeling Method.
IV-F2 Quantitative Analysis
To further quantify the contribution of the EAM, our work conducted a comparative analysis under identical training conditions, evaluating model performance with and without the module on key metrics. The results are presented in Table III.
TABLE III: Impact of the Edge Attention Mechanism on Model Performance (Ablation Study)
EAM ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
92.1% 72.9% 0.701
95.2% 81.6% 0.412
As shown in Table III, integrating the Edge Attention Mechanism significantly improves both ACC and mIoU while reducing the final loss. These results demonstrate that the EAM effectively enhances local boundary precision, enabling the generation of more stable and accurate segmentation labels.
IV-G Efficiency Evaluation
To evaluate computational efficiency, we compare the training time per epoch across different deep learning segmentation models. We conduct all experiments on an NVIDIA 4090D GPU using a training dataset of 3,101 images. The results are shown in Table IV.
TABLE IV: Comparison of Single Epoch Training Time Across Different Segmentation Models (Efficiency Evaluation)
Model Epoch Time (s)
DeepLabv3+ ResNet50 48
DeepLabv3+ ResNet101 65
DeepLabv3+ MobileNet 29
Ours 26
Experimental results indicate that our model achieves a significantly lower training time per epoch compared to other models. Specifically, it reduces training time by approximately 45.8% compared to DeepLabv3+ ResNet50 and 60% compared to DeepLabv3+ ResNet101. These findings highlight the superior training efficiency of our model, enabling faster convergence and lower computational costs.
IV-H Labeling Experiment Analysis
The model’s decoder output classifies each pixel to distinguish between the ocean and oil spills, producing a single-channel grayscale image. Each pixel value represents the probability of belonging to either the ocean or an oil spill. Unlike directly generating binary labels, this output consists of continuous grayscale values ranging from 0 to 255, providing more detailed confidence information.
During the annotation process, threshold-based segmentation inevitably leads to the loss of details, particularly affecting boundary precision and morphological accuracy. As shown in Fig.9, the proposed labeling module effectively preserves more details, especially in the edge contours of oil spill regions. Experimental results further demonstrate that directly using grayscale probability maps allows for a more refined representation of confidence levels across different regions. This approach retains richer information for subsequent analysis and decision-making, making it potentially more advantageous in scenarios requiring human-assisted judgment.
V Conclusion
Our work proposes an edge-enhanced labeling network, called SAM2-ELNet, which combines a labeling module and an edge attention mechanism. This model not only effectively solves the problems of label detail loss, fragmentation, and inaccurate boundaries in manually annotated data, but also has high training efficiency and strong generalization ability. To evaluate the quality of the generated annotations, we conducted experiments on the manually annotated Deep-SAR Oil Spill (SOS) dataset and compared the performance of DeepLab series models trained with original and enhanced labels. The experimental results show that under the same experimental conditions, the model trained with enhanced labels achieves about a 5% improvement in mIoU and accuracy, while also showing faster convergence and a lower final loss value, which indicates that the improved annotations better approximate the real data distribution. For cross-dataset transfer, the model shows strong generalization ability and is able to generate high-fidelity pseudo labels for new datasets. This approach provides a feasible solution for efficient processing of large-scale remote sensing images.
Experiments on synthetic aperture radar (SAR) imagery confirm the model’s effectiveness. Future research will expand its application to multispectral and hyperspectral remote sensing data to assess its robustness across different modalities. Developing a universal segmentation model for remote sensing remains challenging. This study inspires further optimizations in adapter modules and decoder structures to enhance transfer learning, supporting more efficient and accurate image analysis in remote sensing.

### v2

@@
 Abstract
-Remote sensing image segmentation is crucial for environmental monitoring, disaster assessment, and resource management, directly affecting the accuracy and efficiency of surface information extraction.
-The performance of existing supervised models in remote sensing image segmentation tasks highly depends on the quality of label data.
-However, current label data mainly relies on manual annotation, which comes with high time costs and is subject to subjective interference, resulting in distortion of label boundaries and often a loss of detail.
-To solve the above problems, our work proposes an Edge-enhanced Labeling Network, called SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism.
-This model effectively addresses issues such as label detail loss, fragmentation, and inaccurate boundaries.
-Due to the scarcity of manually annotated remote sensing data, the feature extraction capabilities of traditional neural networks are limited.
-Our method uses the Hiera backbone of the pre-trained self-supervised large model segment anything model 2 (SAM2) as the encoder, achieves high-quality and efficient feature extraction even with small samples by fine-tuning on downstream tasks.
-This study compared the training effects of original and enhanced labels on the manually annotated Deep-SAR Oil Spill (SOS) dataset.
-Results showed that the model trained with enhanced labels performed better and had a lower final loss, indicating closer alignment with the real data distribution.
-Our work also explores the potential of extending the model into an efficient automatic annotation framework through generalization experiments, facilitating large-scale remote sensing image interpretation and intelligent recognition.
-Index Terms:
-Remote Sensing Image Segmentation, Automatic Annotation, Edge Attention Mechanism, Binarization Processing, Label Enhancement.
+Remote sensing image segmentation is crucial for environmental monitoring, disaster assessment, and resource management, but its performance largely depends on the quality of the dataset.
+Although several high-quality datasets are broadly accessible, data scarcity remains for specialized tasks like marine oil spill segmentation.
+Such tasks still rely on manual annotation, which is both time-consuming and influenced by subjective human factors.
+The Segment Anything Model 2 (SAM2) has strong potential as an automatic annotation framework but struggles to perform effectively on heterogeneous, low-contrast remote sensing imagery.
+To address these challenges, we introduce a novel label enhancement and automatic annotation framework, termed SAM2-ELNet (Enhancement and Labeling Network).
+Specifically, we employ the frozen Hiera backbone from the pre-trained SAM2 as the encoder, while fine-tuning the adapter and decoder for different remote sensing tasks.
+Additionally, the proposed framework includes a label quality evaluator for filtering, ensuring the reliability of the generated labels.
+We design a series of experiments targeting resource-limited remote sensing tasks and evaluate our method on two datasets: the Deep-SAR Oil Spill (SOS) dataset with Synthetic Aperture Radar (SAR) imagery, and the CHN6-CUG Road dataset with Very High Resolution (VHR) optical imagery.
+The proposed framework can enhance coarse annotations and generate reliable training data under resource-limited conditions.
+Fine-tuned on only 30% of the training data, it generates automatically labeled data.
+A model trained solely on these achieves slightly lower performance than using the full original annotations, while greatly reducing labeling costs and offering a practical solution for large-scale remote sensing interpretation.
+publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
 I Introduction
-Remote sensing typically refers to the technology of imaging the earth system in a specific electromagnetic spectrum band on an aerospace or aviation platform, so as to obtain multi-faceted characteristic information of the observed object.
-It’s imaging methods [1] include optical imaging, thermal infrared [2], hyperspectral [3], [4], and synthetic aperture radar (SAR) [5].
-Currently, there are already various ways to extract valuable information from the massive amount of data.
+Remote sensing typically refers to the technology of imaging the Earth system in a specific electromagnetic spectrum band on an aerospace or aviation platform to obtain multi-faceted characteristic information of the observed object.
+Its imaging methods [1] include optical imaging, thermal infrared [2], hyperspectral [3], [4], and SAR [5].
+Currently, various ways exist to extract valuable information from the massive amount of data.
 Remote sensing image segmentation has become a focal point of contemporary research due to its capability to delineate target regions and reveal terrestrial spatial structures.
@@
 When applied to simple scenes, these methods often yield satisfactory results.
-But their sensitivity to radiometric noise, occlusive shadows, and topological ambiguities leads to issues, especially when dealing with high-dimensional features and complex spatial relationships.
-Driven by advancements in the field of computer vision (CV), deep learning techniques have become the leading approach for remote sensing image segmentation [49], [50], [51], [52], [53].
+However, their sensitivity to radiometric noise, occlusive shadows, and topological ambiguities leads to issues, especially when dealing with high-dimensional features and complex spatial relationships.
+Driven by advancements in the field of computer vision (CV), deep learning techniques have become the leading approach for remote sensing image segmentation [10], [11], [12], [13], [14], [15].
 Supervised learning methods in this domain rely on large-scale, high-quality annotated datasets to effectively train robust models.
@@
 With the growing adoption of these methods, researchers are seeking better ways to optimize manually annotated data while improving labeling efficiency in remote sensing applications.
-Compared with CV tasks, remote sensing images have more severe noise and scale changes, which makes end-to-end self-supervised methods face greater challenges.
+Compared with CV tasks, remote sensing images exhibit more severe noise and scale variations, posing greater challenges for end-to-end self-supervised methods.
 Wang et al.
@@
 Its application to remote sensing image annotation still requires manual input, and its effectiveness remains limited in scenarios involving small object detection, high-noise images, or abnormal data distributions.
-To enhance the application of SAM in remote sensing image segmentation, Zhang et al.
-[36] introduced the adapter-scale and adapter-feature modules to improve SAM’s encoding capability, making it more suitable for semantic segmentation tasks in remote sensing.
+To enhance the applicability of SAM in remote sensing domain, Luo et al.
+[27] proposed the SAM-RSIS framework, providing important insights for subsequent research.
 Ma et al.
@@
 [29] developed the RSPrompter method, which learns to generate category-specific prompts to enhance SAM’s applicability in remote sensing instance segmentation.
-Despite SAM2 performs well in many tasks, it may still generate segmentation results unrelated to specific categories when manual prompts are absent, which limits its usability in the specific context of remote sensing segmentation.
-To address this limitation, Xiong et al.
-[15] integrated SAM2’s hierarchical feature extraction with U-Net’s spatial attention mechanism and proposed a novel network, SAM2-UNet.
-The architecture combines U-Net’s symmetrical decoder with a lightweight parameter adapter [62], [63], enabling efficient fine-tuning for downstream applications.
-Extending the SAM2-UNet framework, we propose a new model for label refinement and automatic annotation, called edge-enhanced labeling network (SAM2-ELNet).
-This model leverages large-scale self-supervised pretrained models for feature extraction and integrates few-shot learning for fine-tuning on downstream tasks.
-Our work introduce an edge attention mechanism and a labeling module to address issues such as label detail loss, fragmentation, and inaccurate boundaries in manually annotated remote sensing segmentation data, with the potential to be extended into an efficient automatic labeling framework.The main contributions of this study are as follows.
+Although SAM2 performs excellently in multiple tasks, its segmentation results often suffer from bias when applied to the automatic annotation process without manual prompts, and adding manual prompts significantly increases the cost.
+Additionally, these base models are typically trained on natural image data, making them difficult to directly apply to remote sensing data, which often features noise, heterogeneity, and low contrast.
+In practical applications, remote sensing data varies significantly across different scenarios, tasks (such as marine oil spill segmentation and road segmentation), and sensors (such as SAR and VHR).
+The performance of general large models in these diverse scenarios, especially in specific application areas, often falls short of ideal results.
+To address this limitation, from a practical application perspective, we propose a novel label enhancement and automatic annotation framework, termed SAM2-ELNet (Enhancement and Labeling Network).
+Unlike semi-supervised or weakly supervised approaches that aim to improve model accuracy, this work focuses on building a data-centric solution for remote sensing segmentation in resource-constrained scenarios.
+Specifically, we aim to improve data usability and quality through automatic annotation and label enhancement, providing a more reliable foundation for downstream tasks.
+The framework is designed to improve label quality in scenarios with large-scale coarse annotations and to enable high-quality automatic labeling in cases where only limited manual annotations are available.
+By freezing the backbone network of SAM2 as the encoder and fine-tuning the decoder and adapter for downstream tasks, we effectively address the domain adaptation challenge of SAM2 across different remote sensing data.
+Our framework also incorporates a label quality evaluator with error-correction feedback, where the high-quality reliable labels are further used for model fine-tuning, while lower-quality labels are re-labeled.
+Since the backbone is frozen, the framework allows for cost-effective, rapid fine-tuning with small datasets, customized applications in small-scale real-world scenarios.
+The proposed framework supports both “low-cost few-shot annotation + large-scale automatic annotation” and “low-cost coarse annotation + label enhancement”, contributing to the advancement of large-scale remote sensing data processing and intelligent interpretation.
+The main contributions of this study are as follows:
   1. 1.
-    To address the issue of suboptimal annotations caused by heterogeneity and noise interference in remote sensing segmentation tasks, our work propose SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism.
-This model effectively optimizes issues such as label detail loss, fragmentation, and inaccurate boundaries.
+    To address the high data annotation costs and inconsistent quality in resource-constrained remote sensing applications, we propose a novel label enhancement and automatic annotation framework, SAM2-ELNet.
+This framework enables label enhancement for existing annotations and automatic annotation for unlabeled data under low-cost conditions.
   2. 2.
-    To enhance the applicability and training efficiency of SAM for specific remote sensing image segmentation tasks, we leverage a pre-trained large model and adopt a fine-tuning strategy with a frozen encoder.
-This approach retains the robust feature extraction capability of SAM2 while significantly improving training efficiency.
+    To generate more reliable labels, we design a label quality evaluator, enabling our framework to incorporate a feedback correction mechanism.
+To enhance the label structure representation of existing models, we adopt the Edge Attention Mechanism (EAM), which strengthens the model’s ability to capture fine boundaries.
   3. 3.
-    In exploring large-scale automatic annotation, we discuss the model’s generalization ability and investigate its potential to be extended into an efficient automatic labeling framework, facilitating the construction of scalable, high-quality datasets.
+    Starting with the question “Can automatic annotation replace manual annotation?”, we developed label quality and automatic annotation evaluation experiments tailored to remote sensing, offering insights for future research.
 The remainder of this paper is organized as follows.
 Section II reviews related works on remote sensing image segmentation, automatic annotation methods, and the application of pretrained models in remote sensing.
-Section III presents the proposed SAM2-ELNet architecture, providing a detailed description of its components.
+Section III presents the proposed SAM2-ELNet framework, providing a detailed description of its components.
 Section IV offers a comprehensive experimental analysis, including model performance evaluation, ablation studies, generalization capability assessment, and efficiency comparisons.
-Both qualitative and quantitative analyses are conducted to validate the effectiveness of the proposed model.
+Both qualitative and quantitative analyses are conducted to validate the effectiveness of the proposed framework.
 Finally, Section V concludes the study.
@@
 In recent years, researchers have actively explored efficient annotation and training strategies to reduce dependence on high-quality manual annotations, lower costs, and enhance the accuracy of remote sensing image segmentation.
+Figure 1: Architecture of SAM2-ELNet.
+The proposed SAM2-ELNet consists of a label generation network and a label quality evaluator.
+It enables label enhancement and automatic annotation in resource-constrained remote sensing segmentation scenarios through fine-tuning.
 Yang et al.[30] introduced the EasySeg framework, which addresses domain adaptation challenges in remote sensing semantic segmentation.
@@
 By comparing noisy label pretraining with self-supervised learning methods such as DINO and MoCo, they further demonstrated the effectiveness of noisy labels in pretraining and examined how class inconsistency influences model transferability.
+Yuan et al.
+[32] combined MLLM with SAM to achieve more flexible and powerful segmentation functionality.
 Li et al.[33] developed the SAM-OIL framework, which integrates YOLOv8-based object detection, an adapted SAM segmentation module, and an ordered mask fusion (OMF) mechanism for oil spill detection in SAR imagery.
@@
 To further enhance annotation quality and efficiency, Wu et al.[34] pioneered the use of vision-language models (VLM) in farmland segmentation.
-They constructed a farmland image-text paired dataset and integrated semantic segmentation models with large language models (LLMs), developing the FSVLM model, which improved generalization and segmentation accuracy of remote sensing segmentation .
+They constructed a farmland image-text paired dataset and integrated semantic segmentation models with large language models (LLMs), developing the FSVLM model, which improved generalization and accuracy of remote sensing segmentation .
 Li et al.[35] introduced a semi-automatic annotation method that combines anchor real-time lines and ROI real-time line algorithms to address label incompleteness and deficiencies in traditional remote sensing indices and interactive segmentation.
@@
 Ma et al.[42] explored geophysical foundation models (GeoFM) for geophysical exploration and applied SAM to seismic phase picking.
-SAM also enhances building segmentation and urban area delineation in remote sensing imagery [30], [31], [32], [33], [34], demonstrating its adaptability in processing complex geographical features.
-Figure 1: Architecture of SAM2-ELNet.
-The proposed SAM2-ELNet comprises an encoder, decoder, receptive field blocks (RFB), and a labeling module.
-It integrates an edge attention mechanism and parameter-efficient fine-tuning (PEFT) to enhance segmentation accuracy and improve label quality for remote sensing imagery.
+SAM also enhances building segmentation and urban area delineation in remote sensing imagery [43], [44], [45], [46] demonstrating its adaptability in processing complex geographical features.
+Figure 2: Architecture of the label generation network.
+The proposed network consists of multiple encoders, decoders, and receptive field blocks (RFB).
+It integrates EAM and parameter-efficient fine-tuning (PEFT) to enhance segmentation accuracy and improve label quality in remote sensing imagery.
+Figure 3: Architecture of the Hiera. The encoder leverages the Hiera backbone for multi-scale feature modeling, with adapters enhancing feature adaptation.
 III Proposed Method
-Building upon the research advancements of SAM2-Unet [15], this paper proposes a model SAM2-ELNet designed for enhanced labeling and automatic annotation.
-This model is specifically tailored for annotation enhancement and automatic labeling, integrating adapters to achieve parameter-efficient fine-tuning (PEFT) while preserving the feature extraction capability of the pre-trained encoder.
+In this section, we introduce the basic framework of SAM2-ELNet.
+The proposed framework is designed for dataset construction and optimization in resource-constrained remote sensing segmentation scenarios.
+It revolves around two core modes: (1) low-cost few-shot annotation + large-scale automatic annotation, and (2) low-cost coarse annotation + label enhancement.
+As shown in Fig.
+1, the framework mainly consists of two modules: a label generation network and a label quality evaluator for label filtering.
+The following subsections provide a detailed description of each module.
+III-A Label Generation Network
+The label generation network is a pre-trained foundation model that can be rapidly fine-tuned for various practical remote sensing segmentation applications, particularly in resource-constrained scenarios.
+Building upon the advancements of SAM2-UNet [47], this paper proposes a label generation network specifically designed for annotation enhancement and automatic labeling.
+By integrating lightweight adapters, the network enables parameter-efficient fine-tuning (PEFT) while retaining the powerful feature extraction capabilities of the original pre-trained encoder.
 As shown in Fig.2, since remote sensing datasets are limited, retraining Hiera from scratch could weaken its feature extraction ability.
@@
 To enable efficient fine-tuning while maintaining adaptability to remote sensing tasks, SAM2-ELNet inherits the adapter design from SAM2-UNet.
-We integrate adapters before the multi-scale modules of the Hiera encoder, allowing the model to dynamically recalibrate feature distributions and reducing the number of tunable parameters.
+We integrate adapters [48], [49] before the multi-scale modules of the Hiera encoder, allowing the model to dynamically recalibrate feature distributions and reducing the number of tunable parameters.
 SAM2-ELNet employs a frozen encoder strategy, its PEFT approach is similar to low-rank adaptation [50] (LoRA).
-LoRA mainly targets the attention mechanism of Transformer structures, but SAM2-ELNet embeds its adapter module before the multi-scale modules of the Hiera encoder to adjust feature mapping and enhance model adaptability, making it more suitable for remote sensing label optimization tasks.
-In the decoder stage, SAM2-ELNet incorporates an edge attention mechanism to sharpen feature responses along object boundaries, improving segmentation accuracy.
-We also introduce a labeling module that transforms continuous grayscale maps into high-fidelity binary labels, preserving critical details.
-The subsequent sections will delve deeper into each module.
-III-A Encoder, Adapter, and RFB Modules
+LoRA mainly targets the attention mechanism of Transformer structures, but SAM2-ELNet embeds its adapter module before the multi-scale modules of the Hiera encoder to adjust feature mapping and enhance model adaptability, making it more suitable for remote sensing label generation tasks.
 Hiera (hierarchical transformer) [51] is a Transformer backbone network designed for efficient visual feature extraction.
 Unlike traditional convolutional neural networks (CNNs), Hiera captures both local details and global semantics through hierarchical feature modeling.
-Its layered computation strategy reduces computational cost while maintaining strong representational power.
-Our model uses the SAM2-pretrained Hiera as an encoder, utilizing its multi-scale feature learning to generate high-quality segmentation features.
-Four receptive field blocks (RFBs) [55], [56] further process the extracted features, reducing the channel count to 64 to enhance lightweight feature representation.The structure of the Hiera is illustrated in Fig.
-2.
+Its layered computation strategy reduces computational cost while maintaining strong representational capacity.
+Our model employs the SAM2-pretrained Hiera as encoder, utilizing its multi-scale feature learning to generate high-quality segmentation features.
+Four receptive field blocks (RFBs) [52], [53] further process the extracted features, reducing the channel count to 64 to enhance lightweight feature representation.
+The structure of the Hiera is illustrated in Fig.
+3.
+Figure 4: Structure of the label quality evaluator.
+The module takes three perturbed inputs processed by differently fine-tuned label generation networks and evaluates prediction consistency.
+Low-quality labels are identified and flagged for re-annotation.
 Hiera’s large parameter count makes full fine-tuning computationally expensive.
@@
 This structure allows the model to efficiently recalibrate feature distributions while reducing the number of trainable parameters.
-III-B Decoder
-The original decoder follows the classical U-Net design, consisting of three decoder blocks, each comprising two Conv-BN-ReLU layers.
-When processing remote sensing images with blurred boundaries or low contrast, this structure tends to produce segmentation outputs with unclear contours, fragmented edges, or excessive smoothing.
-To address these issues, we introduce the edge attention mechanism (EAM) to enhance segmentation precision, particularly in boundary regions.
-Figure 2: Architecture of the Hiera. Proposed in SAM2-Unet, the encoder leverages the Hiera backbone for multi-scale feature modeling, with adapters enhancing feature adaptation.
-EAM enhances the feature map’s focus on boundary regions during the upsampling stage.
-After each upsampling step, it ensures that multi-scale feature aggregation preserves critical boundary information.
-The EAM in this study employs convolution operations to extract object boundary features and applies element-wise weighting to refine the input features, guiding the model to emphasize boundary details.
-The edge attention weight is computed as follows:
+In remote sensing image segmentation tasks, accurately delineating object boundaries plays a critical role, especially when the target objects such as roads, marine oil spills, or man-made structures appear narrow, elongated, or fragmented.
+Inaccurate boundary handling can lead to serious consequences.
+For instance, boundary errors in road segmentation may disrupt connectivity analysis, affecting downstream applications like navigation or path planning.
+In oil spill detection, boundaries often lack clarity and exhibit smooth intensity transitions, making them hard to distinguish from the background.
+Without adequate boundary awareness, models tend to produce overly expanded prediction regions, significantly overestimating the affected area and misleading environmental assessments or emergency responses.
+The baseline decoder follows the classical U-Net design, consisting of three decoding modules, each with two Conv-BN-ReLU layers.
+However, this structure struggles with blurred boundaries or low-contrast regions.
+To address this issue and enhance the model’s ability to capture fine boundary details, we introduce the EAM.
+Unlike traditional convolutional modules that aggregate spatial features uniformly, EAM focuses on the transition zones between foreground and background.
+It strengthens feature responses in edge regions and improves boundary awareness.
+The network applies EAM after each upsampling step to ensure that multi-scale feature fusion retains essential boundary cues.
+Specifically, EAM uses convolutional layers to extract edge-related features and applies element-wise weighting to refine the input features, guiding the model to emphasize edge information.
+The attention weights for edge refinement are computed as follows:
 $A_{E}=\sigma(\text{BN}(W_{\text{conv}}*X))$ (1)
@@
 $X^{\prime}$ is an enhanced feature map emphasizing edge-related information.
-This module adopts a lightweight design, using only a independent 3×3 convolutional layer, which significantly reduces computational cost compared to self-attention mechanisms.
-III-C Loss Function
-We formulate the loss function as a combination of the weighted Intersection over Union (IoU) loss $L_{\text{wIoU}}$ and the weighted binary cross-entropy (BCE) loss $L_{\text{wBCE}}$, defined as:
-$L=L_{\text{wIoU}}+L_{\text{wBCE}}$ (3)
-To enhance training effectiveness, we apply deep supervision to all segmentation outputs $S_{i}$, leading to the final total loss function:
-$L_{\text{total}}=\sum_{i=1}^{3}L(G,S_{i})$ (4)
-where $G$ denotes the ground truth labels, and $S_{i}$ represents the multi-level segmentation outputs generated by the model.
-Figure 3: Labeling Module Architecture. Converts grayscale predictions into binarized labels using edge detection, morphological processing, and the watershed algorithm for precise region extraction.
-III-D Labeling Module
-The model’s initial predictions are grayscale images with intensity values ranging from 0 to 255.
-To transform these predictions into binarized labels, our work design a labeling module based on edge detection and the watershed transform.
-This approach integrates Canny edge detection [41], morphological processing [42], [43], distance transform, and the watershed algorithm [44], [45] to accurately extract foreground, background, and boundary regions.
-The structure of the labeling process is illustrated in Fig.
-3.
-To detail the labeling procedure, we apply the Canny operator to an input image $I$ to extract object contours, emphasizing prominent edges. This process produces the edge map $E$, defined as:
-$E=\text{Canny}(I,T_{\min},T_{\max})$ (5)
-where $E$ represents the binary edge map, while $T_{\min}$ and $T_{\max}$ denote the lower and upper threshold values of the Canny operator.
-This step enhances object boundaries by highlighting regions with sharp intensity transitions.
-To fill gaps within foreground objects, we apply a combination of binarization and morphological dilation. After thresholding the input image, we perform dilation:
-$T(x,y)=\begin{cases}255,&I(x,y)<T_{\text{thresh}}\\ 0,&\text{otherwise}\end{cases}$ (6)
-$D=\text{dilate}(T,K,n)$ (7)
-where $T_{\text{thresh}}$ is the binarization threshold, $K$ is the morphological kernel, and $n$ represents the number of dilation iterations.
-This operation enhances the foreground region, refines boundary continuity, and reduces discontinuities within the object.
-To generate the segmentation labels, we construct the foreground mask $F_{\text{fg}}$ by merging the dilation result $D$ with the Canny edge map $E$, Similarly, we obtain the background mask $F_{\text{bg}}$ by inverting the binarization result:
-$F_{\text{fg}}=D\cup E,\quad F_{\text{bg}}=\neg T$ (8)
-Next, we optimize the foreground segmentation by constraining it with the Euclidean distance transform:
-$D_{T}=\text{distanceTransform}(F_{\text{fg}},L_{2},5)$ (9)
-$F_{\text{fg}}^{\prime}=\begin{cases}255,&D_{T}(x,y)>\alpha\cdot\max(D_{T})\\ 0,&\text{otherwise}\end{cases}$ (10)
-where $\alpha$ is the threshold scaling factor, which controls the extent of the foreground region to ensure boundary precision without excessive expansion.
-Using the extracted foreground and background regions, the labeling matrix $M(x,y)$ is constructed as:
-$M(x,y)=\begin{cases}2,&(x,y)\in F_{\text{fg}}^{\prime}\\ 1,&(x,y)\in F_{\text{bg}}\end{cases}$ (11)
-This matrix serves as the marker input for the watershed transform, which segments the image into foreground, background, and boundary regions based on gradient-based region growing:
-$M^{\prime}=\text{Watershed}(I,M)$ (12)
-The final labeled output $M^{\prime}$ assigns 0 for boundary regions and background, sets 255 for foreground objects, thereby generating high-quality segmentation labels.
-III-E Overview of the Label Enhancement Process
-The overall workflow of the proposed label enhancement method consists of the following stages:
+This module adopts a lightweight design, using only an independent 3×3 convolutional layer, which significantly reduces computational cost compared to self-attention mechanisms.
+III-B Loss Function
+We formulate the loss function as a combination of the weighted Intersection over Union (IoU) loss $L_{\text{wIoU}}$ and the weighted binary cross-entropy (BCE) loss $L_{\text{wBCE}}$, controlled by a balancing factor $\lambda$, defined as:
+$L=\lambda\cdot L_{\text{wIoU}}+(1-\lambda)\cdot L_{\text{wBCE}}$ (3)
+To enhance training effectiveness, we apply deep supervision to all segmentation outputs $S_{i}$, assigning each output a weight $\alpha_{i}$. The final total loss function is defined as:
+$L_{\text{total}}=\sum_{i=1}^{3}\alpha_{i}\cdot L(G,S_{i})$ (4)
+where $L(G,S_{i})$ denotes the loss between the ground truth $G$ and the predicted segmentation output $S_{i}$, and $\alpha_{i}$ is the supervision weight assigned to each level $i$ of the segmentation output.
+The weights $\alpha_{i}$ are normalized such that $\sum\alpha_{i}=1$.
+III-C Label Quality Evaluator
+Automatically generated labels inevitably contain noise and bias in downstream applications.
+To ensure label reliability, this study introduces a label quality evaluator that assesses the consistency of predicted labels and filters out unstable or low-confidence results.
+The structure of the evaluator is illustrated in Fig.
+4.
+Each remote sensing image $I$ undergoes three types of basic data augmentation, such as Gaussian noise, horizontal flipping, and random rotation, resulting in three perturbed versions denoted as $I_{1},I_{2},I_{3}$.
+A label generation network with different fine-tuning epochs processes the three augmented inputs separately, producing segmentation predictions $P_{1},P_{2},P_{3}$.
+The three outputs undergo a mutual alignment process to ensure that all predictions share the same spatial orientation, enabling consistent pixel-wise comparison.
+To quantify pixel-level similarity among the binarized predictions, the root mean squared error (RMSE) is used as the consistency metric.
+The RMSE score $R_{i}$ at each pixel location $i$ is calculated as:
+$R_{i}=\sqrt{\frac{1}{3}\sum_{j=1}^{3}\left(p_{i}^{(j)}-\bar{p}_{i}\right)^{2}}$ (5)
+In this formulation, $p_{i}^{(j)}\in\{0,1\}$ denotes the binary foreground prediction at pixel $i$ in the $j$-th perturbed input. The mean prediction $\bar{p}_{i}$ is computed as:
+$\bar{p}_{i}=\frac{1}{3}\sum_{j=1}^{3}p_{i}^{(j)}$ (6)
+The RMSE reflects the level of disagreement at pixel $i$.
+A lower $R_{i}$ indicates stronger agreement among the three predictions, suggesting a more reliable label.
+To assess the overall consistency of a label, the final RMSE score is obtained by averaging $R_{i}$ over all pixel positions.
+Global consistency across binary segmentation masks is measured using the Intersection over Union (IoU) and Dice coefficient. For any two masks $A$ and $B$, these metrics are computed as:
+$\text{IoU}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (7)
+$\text{Dice}(A,B)=\frac{2|A\cap B|}{|A|+|B|}$ (8)
+Average IoU and Dice scores across all three pairs of predictions $P_{1},P_{2},P_{3}$ are calculated using the following expressions:
+$\text{IoU}_{\text{avg}}=\frac{1}{3}\sum_{1\leq i<j\leq 3}\text{IoU}(P_{i},P_{j})$ (9)
+$\text{Dice}_{\text{avg}}=\frac{1}{3}\sum_{1\leq i<j\leq 3}\text{Dice}(P_{i},P_{j})$ (10)
+A label is considered low-quality if the pixel consistency score $C_{i}$, the average IoU, or the Dice score remains below a predefined threshold.
+To obtain an overall label quality score, the pixel-wise RMSE score $R$, average IoU, and average Dice coefficient are combined using a weighted formulation:
+$Q=\beta_{1}\cdot(1-R)+\beta_{2}\cdot\text{IoU}_{\text{avg}}+\beta_{3}\cdot\text{Dice}_{\text{avg}}$ (11)
+The weighting coefficients $\beta_{1},\beta_{2},\beta_{3}$ reflect the relative contribution of each metric.
+The RMSE term is transformed into 1R1-R to ensure that a higher score consistently indicates better agreement among the predictions.
+A label is marked as low-quality and scheduled for re-annotation if the final quality score $Q$ falls below a predefined threshold.
+III-D Overview of the Label Enhancement and Automatic Annotation Process
+The proposed label enhancement and automatic annotation process consists of the following stages:
   1. 1.
-    Initial Manual Annotation.
-A small subset of manually annotated data serves as the initial training set.
-Although these annotations may exhibit suboptimal quality, they still provide essential supervisory signals.
-This dataset supports the preliminary training of the model, enabling it to learn fundamental segmentation patterns and adapt to the specific characteristics of the target task.
+    Initial Label Preparation.
+Prepare a small set of high-quality manually annotated samples when performing automatic annotation, or construct a large set of coarse manual annotations when conducting label enhancement.
+These datasets serve as the initial supervision sources for model fine-tuning.
   2. 2.
     Model Fine-Tuning.
-The model initializes with pretrained weights and fine-tunes on the manually annotated dataset.
-With the encoder weights remaining frozen, the fine-tuning process primarily optimizes the adapters and decoder.
-This strategy mitigates the adverse impact of suboptimal quality annotations on model learning, facilitating better adaptation to remote sensing segmentation tasks.
+Fine-tune the model using the prepared initial labels.
+The model initializes with pretrained weights, while only the adapters and decoder are updated during training to preserve the generalization ability of the encoder.
   3. 3.
-    Label Generation. Upon completion of training, the fine-tuned model infers segmentation results, automatically generating high-quality segmentation labels.
+    Automatic Label Generation and Filtering.
+Apply the fine-tuned model to generate segmentation labels.
+The label quality evaluator filters out low-quality predictions, which are flagged for re-annotation to ensure label reliability.
   4. 4.
-    Iterative Optimization (Optional).
-The model can further fine-tune using the enhanced labels, iteratively refining segmentation quality through adaptive learning.
-This iterative strategy progressively improves annotation accuracy and stability.
+    Model Refinement with Filtered Labels.
+Incorporate the retained high-quality labels into the training set and further fine-tune the model.
+This refined model can generate more accurate annotations in subsequent iterations.
+  5. 5.
+    Iterative Enhancement.
+Repeat the label generation, quality filtering, and model refinement steps iteratively.
+This progressive process incrementally improves annotation quality and model performance in resource-constrained scenarios.
 IV Experiments and Discussion
 IV-A Datasets
+IV-A1 Deep-SAR Oil Spill (SOS) dataset
 The Deep-SAR Oil Spill (SOS) dataset consists of SAR images from two sources [55]: PALSAR data from the Mexican oil spill region and Sentinel data from the Persian Gulf.
@@
 Illustrative examples are shown in Fig.
-4.
+5.
 (a) Test sample 10011
@@
 (d) The annotation misclassifies a sea region as an oil spill and exhibits irregular polygonal boundaries.
+IV-A2 CHN6-CUG Road dataset
+The CHN6-CUG Road dataset [56] is a high-resolution optical road segmentation dataset constructed from Google Earth imagery.
+It covers six representative cities in China with varying urban characteristics: Beijing (Chaoyang), Shanghai (Yangpu), Wuhan, Shenzhen (Nanshan), Hong Kong (Sha Tin), and Macao.
+The dataset includes 4,511 manually annotated images with a resolution of 50 cm/pixel and size $512\times 512$, of which 3,608 are used for training and 903 for testing.
+Labeled roads encompass various transportation types, including railways, highways, urban, and rural roads.
+Due to occlusions and complex urban structures, the annotations may exhibit challenges such as partial coverage and blurred boundaries, making this dataset suitable for evaluating segmentation robustness under realistic conditions.
 IV-B Hyperparameter Settings
@@
 IV-D1 Qualitative Analysis
-The model fine-tunes using only the training subset of the PALSAR dataset for 200 epochs.
-After training, the model enhances both the training and test set labels of the PALSAR dataset, generating high-quality annotations for subsequent analysis.
-From a qualitative perspective, the enhanced labels show substantial visual improvements over the original annotations, as illustrated in Fig. 5.
+To simulate the practical application scenario of “low-cost coarse annotation + label enhancement” we treat the label annotations in the Deep-SAR Oil Spill (SOS) dataset as coarse labels to be enhanced.
+Both quantitative and qualitative evaluations are conducted accordingly.
+The model is fine-tuned using only the training subset of the PALSAR dataset for 200 epochs.
+After training, it enhances both the training and test set labels of the PALSAR dataset, generating high-quality annotations for subsequent analysis.
+As shown in Fig.
+6, the enhanced labels exhibit smoother and more continuous boundaries, and significantly refine annotation details compared to the original labels.
+This improvement appears consistently across the dataset.
+The enhanced annotations provide better object shape precision and more accurate representation of complex regions and ambiguous edges, enabling clearer delineation of target contours.
+These results demonstrate the model’s ability to effectively optimize manual annotations.
 (a) original image
-(b) manual annotation
-(c) enhanced label
+(b) manual
+(c) U-Net
+(d) SAM2(Guidance)
+(e) SAM2-UNet
+(f) LOGCAN++
+(g) Ours(enhanced)
+($a$)
+Figure 6: Comparison of label enhancement results across different methods.
+Each group shows the original SAR image, the manual annotation, and enhanced labels generated by our model, U-Net (same pipeline), SAM2 (guided by manual labels), SAM2-UNet, and LOGCAN++.
+Our method produces labels with finer detail, more accurate boundaries, and overall higher annotation quality, achieving comparable performance to current state-of-the-art approaches.
+To further evaluate the model, we trained a U-Net using the same pipeline.
+U-Net [57] showed moderate improvements, but our model produced more precise and consistent annotations.
+We also evaluated SAM2 [25] in two variants: SAM2 (guided by manual labels) and SAM2-UNet.
+Compared to SAM2 (guided by manual labels), our model achieved noticeably better boundaries and structure.
+Compared to SAM2-UNet, it provided slightly finer details and more stable quality.
+In addition, we compared our approach with the current SOTA segmentation model LOGCAN++ [58], and achieved comparable performance, further demonstrating the effectiveness of our method in label quality enhancement tasks.
+IV-D2 Quantitative Analysis
+Our objective is to generate higher-quality annotations.
+Directly evaluating models using only the original labels may fail to fully reflect the effectiveness of the enhanced annotations.
+Moreover, since segmentation models tend to exhibit low sensitivity to label quality during training, the presence of noisy annotations may even improve model robustness.
+To address these limitations, we design a new evaluation strategy that constructs multiple test sets to more accurately assess annotation quality.
+TABLE I: Label Quality Evaluation Using Third-Party Segmentation Models
+Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
+DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
+Test-Orig 0.7541 0.0274 0.9157 0.0127
+Test-Enh 0.7877 0.0062 0.9328 0.0045
+U-Net Original Test-HQ 0.8276 - 0.9462 -
+Test-Orig 0.7919 0.0357 0.9350 0.0112
+Test-Enh 0.8369 0.0093 0.9491 0.0029
+TABLE II: Evaluation of Enhanced Labels Using Baseline Models
+Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
+DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
+Test-U-Net 0.8517 0.0701 0.9466 0.0182
+Test-SAM2-UNet 0.8192 0.0377 0.9360 0.0076
+Test-LOGCAN++ 0.7893 0.0077 0.9324 0.0041
+Test-Ours 0.7877 0.0062 0.9328 0.0045
+Specifically, a group of domain experts were invited to refine the annotations of selected samples in the test set, forming a high-quality benchmark test set referred to as Test-HQ.
+Corresponding samples from the original and enhanced labels were also selected to construct two comparison test sets, Test-Orig and Test-Enh.
+A third-party segmentation model undergoes training with the original training set and evaluation on the three test sets.
+Performance metrics offer an indirect means to evaluate the actual quality of both original and enhanced labels.
+Although traditional metrics may not fully reflect the quality of annotations, the performance gap between these results and Test-HQ enables a more objective assessment of label quality.
+(a) Loss
+(b) mIoU
+(c) Accuracy
+Figure 7: Evaluation results of the model trained on original labels across three different test sets.
+The performance on Test-Enh is notably closer to that on Test-HQ, indicating the higher quality of the enhanced annotations.
+(a) original image
+(b) manual labels
+(c) auto-labels
 (d) original image
-(e) manual annotation
-(f) enhanced label
-Figure 5: Results of the label enhancement experiment.
-Each triplet consists of an original SAR image, a manually annotated label, and a model-enhanced label.
-The enhanced labels exhibit smoother and more continuous boundaries, improved annotation details, and refined structural consistency.
-As depicted in Fig.
-5, the enhanced labels exhibit smoother and more continuous boundaries, significantly refining annotation details compared to the original labels.
-This improvement is not limited to isolated cases but is a consistent trend across the dataset.
-The enhanced labels demonstrate superior performance in boundary smoothness, object shape precision, and the accurate depiction of complex regions and ambiguous edges, enabling a more precise delineation of target contours.
-The increased consistency and accuracy of the annotations validate the model’s ability to effectively optimize manually labeled data, offering a more reliable supervisory dataset for remote sensing image segmentation tasks.
-IV-D2 Quantitative Analysis
-In this experiment, we conduct comparative studies using the open-source implementation of DeepLabv3+ [46], [47] to analyze the impact of annotation quality on model performance.
-Specifically, we compare the performance of models trained on original labels versus those trained on enhanced labels.
-All experiments follow identical training conditions, utilizing DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as baseline models.
-Each model is initialized with the same pretrained weights and trained with a consistent set of hyperparameters for 20 epochs.
-TABLE I: Impact of Different Label Qualities on Model Performance (DeepLabv3+ Training Comparison)
-Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
-MobileNet Original Label 0.9251 0.7436 0.0773
-Enhanced Label 0.9523 0.8164 0.0452
-ResNet50 Original Label 0.9195 0.7396 0.0815
-Enhanced Label 0.9476 0.8092 0.0458
-ResNet101 Original Label 0.9210 0.7652 0.0792
-Enhanced Label 0.9337 0.8136 0.0453
-Experimental results Table I and Fig.
-6 indicate that models trained with enhanced labels consistently outperform those trained with original labels in terms of mean Intersection over Union (mIoU) and overall accuracy (ACC).
-This demonstrates that enhanced labels facilitate more effective learning of representative features.
-Furthermore, models trained with enhanced labels converge faster and achieve lower final loss values, confirming that enhanced labels more accurately capture the true data distribution.
-As a result, leveraging enhanced labels significantly enhances overall performance in segmentation tasks.
-(a) overall loss curve
-(b) detailed loss segment
-(c) overall accuracy curve
-(d) overall mIoU curve
-(e) initial loss curve
-Figure 6: Training Convergence and Performance Metrics of DeepLabV3+ ResNet50 on Original and Enhanced Labels.
-(Loss, accuracy (ACC), and mean Intersection over Union (mIoU) trends during training on original and enhanced labels.
-Each epoch consists of 190 iterations, resulting in a total of 3,800 iterations over 20 epochs.)
-IV-E Generalization Performance Evaluation
-To evaluate the model’s generalization ability, we fine-tuned it solely on the PALSAR dataset’s training set, optimizing it over 200 epochs.
-After training, we used the fine-tuned model to refine both the training and testing labels of the Sentinel dataset, producing higher-quality annotations.
-This approach helped us assess the model’s adaptability to different data distributions and explore its potential for automatic annotation tasks.
+(e) manual labels
+(f) auto-labels
+Figure 8: Generalization performance of automatic annotation on the Sentinel dataset.
+The automatically generated labels not only preserve key target structures but also exhibit improved quality in certain regions compared to the original annotations.
+In this experiment, we adopt open-source implementations of DeepLabv3+ and U-Net [55, 59, 57] to analyze the quality of the generated annotations.
+Specifically, the performance of models trained on original labels is evaluated across the three test sets, and the performance gap relative to Test-HQ is calculated to assess label quality more accurately.
+All experiments are conducted under the same training conditions.
+The backbone networks for DeepLabv3+ include MobileNet, ResNet50, and ResNet101.
+Each model is initialized with the same pretrained weights and trained for 20 epochs using consistent hyperparameters.
+The experimental results in Table I and Fig.
+7 show that models trained on the original labels perform significantly better on both Test-Enh and Test-HQ in terms of mean Intersection over Union (mIoU) and accuracy (ACC), compared to their performance on Test-Orig.
+Moreover, the results on Test-Enh are closely aligned with those on Test-HQ, further demonstrating that the enhanced labels have achieved a quality level comparable to expert-level fine annotations.
+To verify the effectiveness of our method, we applied the same evaluation protocol to several representative baseline models. The results are summarized in Table II.
+When using model-generated labels as test sets, evaluation scores often exceed those obtained with manual annotations.
+This happens because coarse or overly smooth test labels make it easier for the model to produce matching predictions, thus inflating the evaluation metrics.
+Visual inspection reveals that labels generated by U-Net tend to cover broader regions.
+When used as test labels, they yield higher scores and exhibit a larger gap from the high-quality reference set (Test-HQ).
+In contrast, our method consistently delivers results that are closer to Test-HQ, indicating that the generated labels better approximate expert-level fine annotations.
+Furthermore, our approach achieves comparable performance to the state-of-the-art model LOGCAN++, reinforcing the reliability of the proposed evaluation scheme and confirming the effectiveness of the label enhancement framework.
+(a) Loss
+(b) mIoU
+(c) Accuracy
+Figure 9: Comparison of model performance trained on auto-annotated versus original labels.
+The results demonstrate comparable performance, confirming the practical effectiveness of automatic annotation.
+(a) original image
+(b) manual labels
+(c) auto-labels
+(d) original image
+(e) manual labels
+(f) auto-labels
+Figure 10: Visual Comparison of Auto-Annotated and Manually Labeled Data.
+The auto-annotated labels exhibit comparable quality to manual annotations, effectively capturing key structures and fine boundaries in high-resolution remote sensing images.
+IV-E Automatic Annotation Performance Evaluation
+To further evaluate the generalization capability of the proposed model, particularly its potential in automatic annotation tasks, we design a new set of experiments to simulate a typical cross dataset annotation scenario.
+Specifically, the model is fine-tuned on the training set of the PALSAR dataset for 200 epochs, and then used to automatically generate labels for the training set of the Sentinel dataset.
 IV-E1 Qualitative Analysis
-To visually assess the impact of label enhancement on segmentation performance in the Sentinel dataset, we conducted a comparative analysis through visual inspection.
+To visually assess the quality of automatically generated labels in the Sentinel dataset, we conducted a comparative analysis through visual inspection.
 As shown in Fig.
-7, models trained with enhanced labels outperform those trained on original labels, demonstrating sharper boundaries, finer segmentation details, and better preservation of target regions.
-The enhancement process effectively mitigates issues like boundary ambiguity and object omission, resulting in more precise segmentation outcomes.
-(a) original image
-(b) manual annotation
-(c) enhanced label
-(d) original image
-(e) manual annotation
-(f) enhanced label
-Figure 7: Generalization Performance of Label Enhancement on the Sentinel Dataset.
-The enhanced labels improve boundary sharpness and segmentation accuracy, demonstrating the model’s adaptability to different data distributions.
+8, the automatically generated labels demonstrate superior quality compared to the original annotations, characterized by clearer boundaries, finer details, and improved representation of target regions.
 IV-E2 Quantitative Analysis
-To further quantify the impact of label enhancement on the model’s generalization performance, we conducted comparative experiments under identical training conditions.
-Specifically, we used DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as benchmark models.
-We initialized all models with the same pre-trained weights, applied identical hyperparameter configurations, and trained them for 20 epochs to evaluate their adaptability and generalization across different datasets.
-TABLE II: Impact of Different Label Qualities on Model Generalization Performance (Cross-Dataset Transfer Experiment)
-Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
-MobileNet Original Label 0.8347 0.7024 0.1073
-Enhanced Label 0.8941 0.7953 0.0606
-ResNet50 Original Label 0.8702 0.7525 0.0976
-Enhanced Label 0.9056 0.8166 0.0597
-ResNet101 Original Label 0.8491 0.7192 0.0983
-Enhanced Label 0.9100 0.8261 0.0554
-Experimental results in Table II indicate that although overall performance declined after cross-dataset transfer, the models still exhibited strong automatic annotation capabilities, further confirming the feasibility of the proposed approach for automatic annotation in remote sensing imagery.
-IV-F Ablation Study: Effect of the Edge Attention Mechanism
-To investigate the impact of the edge attention mechanism on remote sensing image segmentation performance, our work conducted an ablation study.
+To fairly assess the impact of different labels on model performance, we use the same test set across all comparisons—the original test set of the Sentinel dataset.
+Two models are trained separately: one using the automatically generated labels and the other using the original labels.
+By comparing their performance on the same test set, we can effectively evaluate the quality and practicality of the automatically generated annotations.
+This experiment helps analyze the model’s ability to adapt to distribution shifts and confirms its potential for cross-dataset automatic annotation tasks.
+To ensure consistent evaluation, we further adopt two mainstream semantic segmentation models: DeepLabv3+ (with ResNet50 as the backbone) and U-Net.
+All models are initialized with the same pretrained weights and trained under identical hyperparameter settings for 20 epochs.
+Performance on the shared test set is compared to assess model adaptability in cross-domain scenarios and validate the feasibility and effectiveness of the proposed framework for automatic annotation.
+TABLE III: Performance Comparison Between Models Trained on Manual and Auto-Annotated Labels (Sentinel Test Set)
+Test Set Model Train Set ACC mIoU
+Original DeepLabv3+ Manual Labels 0.8577 0.7387
+DeepLabv3+ Auto Labels 0.8566 0.7352
+Original U-Net Manual Labels 0.8781 0.7672
+U-Net Auto Labels 0.8722 0.7649
+As shown in Table III and Fig.
+8, the models trained with auto-annotated labels achieved performance comparable to those trained with original labels on the Sentinel dataset, further confirming the feasibility and effectiveness of the proposed framework in remote sensing applications.
+IV-F Automatic Annotation on Very High-Resolution Remote Sensing Data
+To further validate the applicability of the proposed automatic annotation framework in optical remote sensing scenarios, we conducted an additional set of experiments.
+Unlike the previous SAR-based experiments, this evaluation focuses on the CHN6-CUG Road Dataset, a high-resolution optical remote sensing dataset.
+This setup allows us to assess the generalization capability and practical value of our method under different sensor types and stricter annotation standards.
+To simulate a realistic “low-cost few-shot annotation + large-scale automatic annotation” workflow, only the first 30% of the training data from the dataset was used to fine-tune the model.
+The fine-tuned model then automatically annotated the remaining training data to generate labels.
+We trained two independent models: one using the original labels and the other using the automatically generated labels, and evaluated their performance on a shared test set.
+This experiment aims to examine whether the proposed framework can efficiently annotate large-scale data with limited high-quality supervision and provide reliable guidance for downstream model training under resource-constrained conditions.
+TABLE IV: Performance Comparison Between Models Trained on Manual and Auto-Annotated Labels (CHN6-CUG Road Dataset)
+Test Set Model Train Set ACC mIoU
+Original DeepLabv3+ Manual Labels 0.9581 0.6865
+DeepLabv3+ Auto Labels 0.9548 0.6561
+Original U-Net Manual Labels 0.9556 0.6548
+U-Net Auto Labels 0.9485 0.6379
+As shown in Table IV and Fig.
+10, models trained on automatically annotated data exhibit slightly lower performance compared to those trained on manual labels.
+However, the results remain practically valuable, offering a cost-effective solution for real-world applications.
+IV-G Application of the Label Quality Evaluator
+To evaluate the effectiveness of the proposed Label Quality Evaluator (LQE), we conducted an additional experiment on the CHN6-CUG Road dataset.
+Building upon earlier experiments that compared manually annotated labels and unfiltered auto-generated labels, this study introduced a third setting—auto-generated labels filtered by LQE.
+As shown in Table V, models trained with unfiltered auto-generated labels exhibited a clear decline in both accuracy (ACC) and mean Intersection over Union (mIoU) compared to those trained with manual annotations.
+This confirms that label noise can negatively affect segmentation performance.
+However, LQE was applied to screen the auto-generated labels, and model performance improved significantly.
+In particular, mIoU showed a notable increase.
+To further explore the influence of the Iterative Enhancement hyperparameter, we also tested different loop numbers in the iterative process.
+As shown in Table VI, we observed that the number of loops affected model performance.
+Considering both performance and time cost, we recommend using 2 or 3 iterations as a balanced choice for optimal results.
+These results validate the utility of the Label Quality Evaluator.
+LQE distinguishes high-quality labels from noisy ones, allowing the model to learn from more reliable supervision.
+This makes LQE a valuable component in scalable annotation pipelines, especially under limited annotation resources.
+TABLE V: Experimental Results Evaluating the Effectiveness of the Label Quality Evaluator (LQE)
+Test Set Model Train Set ACC mIoU
+Original DeepLabv3+ Manual Labels 0.9581 0.6865
+DeepLabv3+ Auto Labels 0.9548 0.6561
+DeepLabv3+ Auto Labels (LQE) 0.9572 0.6705
+Original U-Net Manual Labels 0.9556 0.6548
+U-Net Auto Labels 0.9485 0.6379
+U-Net Auto Labels (LQE) 0.9532 0.6469
+TABLE VI: Experimental Results of Iterative Enhancement with Different Loop Numbers
+Loop Numbers ACC mIoU
+0 0.9548 0.6561
+1 0.9575 0.6632
+2 0.9556 0.6686
+3 0.9572 0.6705
+4 0.9569 0.6697
+IV-H Ablation Study: Effect of the EAM
+To investigate the impact of the EAM on remote sensing image segmentation performance, our work conducted an ablation study.
 By comparing model configurations with and without the EAM, our work evaluated differences in both qualitative visualization and quantitative metrics to assess its contribution to improving segmentation accuracy.
-IV-F1 Qualitative Analysis
-Fig.8 presents the segmentation results obtained with (left) and without (right) the Edge Attention Mechanism.
-Figure 8: Impact of the Edge Attention Mechanism on Segmentation Results.
-(b) Without Edge Attention Module, (c) With Edge Attention Mechanism; (e) Without Edge Attention Mechanism, (f) With Edge Attention Mechanism.
-Experimental results show that the Edge Attention Mechanism significantly improves segmentation accuracy along object boundaries.
+IV-H1 Qualitative Analysis
+Fig. 11 presents the segmentation results obtained with (left) and without (right) the EAM.
+Figure 11: Impact of the EAM on Segmentation Results. (b) Without EAM, (c) With EAM; (e) Without EAM, (f) With EAM.
+Experimental results show that the EAM significantly improves segmentation accuracy along object boundaries.
 In regions with low contrast or blurred edges, this module captures finer details, generating smoother and more continuous segmentation boundaries while reducing jagged or fragmented edges.
@@
 From an annotation perspective, incorporating edge attention is more suitable for generating segmentation labels, especially in scenarios where clear boundary delineation is crucial.
-Figure 9: Comparison of Different Labeling Methods. (a) Original SAR Image, (b) Decoder Output, (c) Threshold-Based Binarization, and (d) Proposed Labeling Method.
-IV-F2 Quantitative Analysis
-To further quantify the contribution of the EAM, our work conducted a comparative analysis under identical training conditions, evaluating model performance with and without the module on key metrics.
-The results are presented in Table III.
-TABLE III: Impact of the Edge Attention Mechanism on Model Performance (Ablation Study)
-EAM ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
-92.1% 72.9% 0.701
-95.2% 81.6% 0.412
-As shown in Table III, integrating the Edge Attention Mechanism significantly improves both ACC and mIoU while reducing the final loss.
-These results demonstrate that the EAM effectively enhances local boundary precision, enabling the generation of more stable and accurate segmentation labels.
-IV-G Efficiency Evaluation
+IV-H2 Quantitative Analysis
+To further evaluate the contribution of EAM, we constructed two different test sets by enabling and disabling EAM, and compared them with the baseline performance on TEST-HQ.
+The results are shown in Table VII.
+TABLE VII: Impact of the EAM on Model Performance (Compared with TEST-HQ Baseline)
+Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
+DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
+Original Test-NoEMA 0.8126 0.0311 0.9391 0.0108
+Original Test-EMA 0.7877 0.0062 0.9328 0.0045
+As shown in Table VII, the test set generated with the EAM produces evaluation results that more closely align with those on the finely annotated high-quality Test-HQ dataset, compared to the version generated without EAM.
+This suggests that EAM contributes to the generation of higher-quality labels.
+By incorporating edge-aware constraints during label generation, the model better preserves fine boundary structures, leading to more precise and reliable automatic annotations.
+IV-I Efficiency Evaluation
 To evaluate computational efficiency, we compare the training time per epoch across different deep learning segmentation models.
 We conduct all experiments on an NVIDIA 4090D GPU using a training dataset of 3,101 images.
-The results are shown in Table IV.
-TABLE IV: Comparison of Single Epoch Training Time Across Different Segmentation Models (Efficiency Evaluation)
+The results are shown in Table VIII.
+TABLE VIII: Comparison of Single Epoch Training Time Across Different Segmentation Models (Efficiency Evaluation)
 Model Epoch Time (s)
@@
 DeepLabv3+ MobileNet 29
+LOGCAN++ 24
 Ours 26
-Experimental results indicate that our model achieves a significantly lower training time per epoch compared to other models.
-Specifically, it reduces training time by approximately 45.8% compared to DeepLabv3+ ResNet50 and 60% compared to DeepLabv3+ ResNet101.
-These findings highlight the superior training efficiency of our model, enabling faster convergence and lower computational costs.
-IV-H Labeling Experiment Analysis
-The model’s decoder output classifies each pixel to distinguish between the ocean and oil spills, producing a single-channel grayscale image.
-Each pixel value represents the probability of belonging to either the ocean or an oil spill.
-Unlike directly generating binary labels, this output consists of continuous grayscale values ranging from 0 to 255, providing more detailed confidence information.
-During the annotation process, threshold-based segmentation inevitably leads to the loss of details, particularly affecting boundary precision and morphological accuracy.
-As shown in Fig.9, the proposed labeling module effectively preserves more details, especially in the edge contours of oil spill regions.
-Experimental results further demonstrate that directly using grayscale probability maps allows for a more refined representation of confidence levels across different regions.
-This approach retains richer information for subsequent analysis and decision-making, making it potentially more advantageous in scenarios requiring human-assisted judgment.
+Experimental results indicate that our model achieves significantly lower training time per epoch compared to other models.
+Specifically, it reduces training time by approximately 45.8% compared to DeepLabv3+ with ResNet50 and 60% compared to DeepLabv3+ with ResNet101.
+Moreover, the training time of our model is comparable to that of lightweight architectures such as LOGCAN++ and DeepLabv3+ with MobileNet, demonstrating that it maintains high computational efficiency while preserving segmentation accuracy.
+IV-J Discussion
+In practical remote sensing segmentation tasks, obtaining large-scale, high-quality pixel-level annotations is often infeasible, especially for new regions, sensors, or task types with scarce labeled data.
+SAM2-ELNet provides an efficient way to generate new training data from limited annotations through automatic labeling and label enhancement.
+It is particularly suitable for the following scenarios:
+  1. 1.
+    Post-processing of coarse annotations.
+Given existing coarse-grained labels, SAM2-ELNet can enhance boundary accuracy and structural completeness, making the labels more suitable for refined model training or evaluation.
+  2. 2.
+    Low-shot annotation-driven automation.
+When only a small amount of annotated data is available at the early stage of a project, the model can be fine-tuned with this data and then used to automatically label the entire image set, rapidly generating full-coverage labels for pretraining or iterative refinement.
+  3. 3.
+    Rapid cross-region deployment.
+When adapting a trained model to new regions or image sources, SAM2-ELNet can be quickly fine-tuned with a small number of local samples, reducing re-annotation costs and accelerating deployment.
+The framework requires minimal computational resources for fine-tuning, and in some cases, its automatically generated labels can partially replace manual annotations.
+This makes it suitable for remote sensing projects with limited resources and high annotation costs.
+For targets with vague boundaries or ambiguous semantics, manual verification may still be necessary.
+Future enhancements to the label quality evaluator could further improve reliability.
 V Conclusion
-Our work proposes an edge-enhanced labeling network, called SAM2-ELNet, which combines a labeling module and an edge attention mechanism.
-This model not only effectively solves the problems of label detail loss, fragmentation, and inaccurate boundaries in manually annotated data, but also has high training efficiency and strong generalization ability.
-To evaluate the quality of the generated annotations, we conducted experiments on the manually annotated Deep-SAR Oil Spill (SOS) dataset and compared the performance of DeepLab series models trained with original and enhanced labels.
-The experimental results show that under the same experimental conditions, the model trained with enhanced labels achieves about a 5% improvement in mIoU and accuracy, while also showing faster convergence and a lower final loss value, which indicates that the improved annotations better approximate the real data distribution.
-For cross-dataset transfer, the model shows strong generalization ability and is able to generate high-fidelity pseudo labels for new datasets.
-This approach provides a feasible solution for efficient processing of large-scale remote sensing images.
-Experiments on synthetic aperture radar (SAR) imagery confirm the model’s effectiveness.
-Future research will expand its application to multispectral and hyperspectral remote sensing data to assess its robustness across different modalities.
-Developing a universal segmentation model for remote sensing remains challenging.
-This study inspires further optimizations in adapter modules and decoder structures to enhance transfer learning, supporting more efficient and accurate image analysis in remote sensing.
+This paper proposes SAM2-ELNet, a label enhancement and automatic annotation framework tailored for remote sensing under resource constraints.
+It freezes the SAM2 backbone and fine-tunes lightweight modules to quickly adapt to diverse segmentation tasks with limited annotations.
+To improve label quality, it incorporates an EAM for boundary construction and a label quality evaluator for filtering and feedback correction.
+Experiments on the Deep-SAR Oil Spill and CHN6-CUG Road datasets demonstrate that SAM2-ELNet achieves both effective label enhancement and reliable automatic annotation.
+With only 30% of the annotated data used for fine-tuning, the model automatically labeled the entire dataset.
+Training on these generated labels yielded slightly lower accuracy than the model trained with the full set of original manual annotations, but still confirmed the practicality of the automatic annotation mechanism.
+Future work includes extending SAM2-ELNet to multi-class and multi-modal annotation by incorporating textual semantics for vision-language fusion.
+The label quality evaluator may also be redesigned as a discriminator in an adversarial framework to enable self-improving learning.
+Research on automatic annotation is expected to provide an efficient and scalable solution for large-scale intelligent interpretation in remote sensing.